# ACES: Automatic Cohort Extraction System for Event-Stream Datasets

### v1

@@ Abstract
Reproducibility remains a significant challenge in machine learning (ML) for healthcare. In this field, datasets, model pipelines, and even task/cohort definitions are often private, leading to a significant barrier in sharing, iterating, and understanding ML results on electronic health record (EHR) datasets. In this paper, we address a significant part of this problem by introducing the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). This tool is designed to simultaneously simplify the development of task/cohorts for ML in healthcare and enable the reproduction of these cohorts, both at an exact level for single datasets and at a conceptual level across datasets. To accomplish this, ACES provides (1) a highly intuitive and expressive configuration language for defining both dataset-specific concepts and dataset-agnostic inclusion/exclusion criteria, and (2) a pipeline to automatically extract patient records that meet these defined criteria from real-world data. ACES can be automatically applied to any dataset in either the Medical Event Data Standard (MEDS) or EventStreamGPT (ESGPT) formats, or to any dataset for which the necessary task-specific predicates can be extracted in an event-stream form. ACES has the potential to significantly lower the barrier to entry for defining ML tasks, redefine the way researchers interact with EHR datasets, and significantly improve the state of reproducibility for ML studies in this modality. ACES is available at https://github.com/justin13601/aces.
@@ 1 Introduction
Machine learning (ML) for healthcare suffers from a severe and systemic reproducibility crisis [9]. This challenge is further exacerbated by the need to maintain private and secure datasets, but even with public datasets, ML pipelines are not reliably reproducible from published papers alone. For instance, in numerous attempts to reproduce ML for healthcare studies using the MIMIC-III dataset [6], Johnson et al. found that more than half the time, the cohorts described in the studies could not be reliably reconstructed [5].
@@
This burden in reproducing the basic task and problem definitions in ML for healthcare studies is profoundly detrimental. Beyond the obvious concerns it raises around the robustness of reported results and their readiness for deployment, our inability to reliably define shared, canonical, reproducible task definitions limits our capacity to perform meaningful model comparisons during methodological development. This is particularly notable in settings where not all researchers have mutual access to all datasets, as is common in healthcare. Given the critical role that open benchmarks play in the advancement of ML methods [23, 17, 19], this deficit directly translates to a significant barrier in our ability, as a research community, to effectively experiment, iterate, and develop new ML methodologies in the healthcare space.
@@
Given the clear import of this problem, the research community has naturally explored a number of prospective solutions. These can be largely categorized into two areas: (1) leveraging existing common data models (CDMs) to define reproducible task cohorts only for datasets within these schemas, and (2) defining static benchmarking tasks on individual public datasets. Both of these areas have generated numerous high-impact works. For example, in the area of CDM-driven tools, systems such as the ATLAS tool [13] for OHDSI’s OMOP CDM [12], i2b2’s query tool for the i2b2 CDM [11], and institution-specific tools such as the Stanford Research Repository (STARR) OMOP system and the NIHR Infections in Oxfordshire Database (IORD) cohort discovery platform have all been used to drive numerous new lines of inquiry. Unfortunately, these tools are also extremely limited in that they can only apply to the specific CDM or institutional data warehouse for which they have been defined. Further, because many of these CDMs have had limited penetrance into healthcare’s high-capacity, deep learning ecosystems, they are particularly ill-posed for task and cohort extraction within the deep learning communities. Conversely, public static benchmarks [3, 8, 20] over datasets such as MIMIC-IV [4] or eICU [15] have also been extremely impactful for the ML community. However, they are all tied to only a single or small set of datasets and tasks. Given the highly dynamic nature of clinical data and healthcare requirements, this is insufficient for the benchmarking and reproducibility needs faced by the ML for healthcare community.
@@
1. The Need for Interoperability The limited public datasets and only partially used CDMs cannot capture the diverse clinical populations, needs, and model capacities necessary for tangible ML progress in healthcare. To address this, systems for automated task extraction must be meaningfully interoperable across both public and private datasets with diverse input schemas.
2. The Need for Flexibility A single, static benchmark cannot encompass the variety of clinical tasks relevant to clinicians and informaticians. As existing tools (with only limited interfaces to define queries of per-set vocabularies) may struggle to generalize to new clinical tasks, ideal solutions must be flexible enough to accommodate a myriad of new task definitions, criteria formats, and disease or deployment areas.
3. The Need for Accessibility, Usability, & Applicability in Deep Learning Workflows While many existing tools present with no-code interfaces (e.g., web tools to build queries) that are essential for less technically literate audiences, integrating such tools with deep learning workflows can prove challenging. Deep learning systems are often run in a semi-programmatic manner on siloed, private computational clusters where researchers have minimal control. Hence, existing tools can be significantly inconvenient. Instead, successful tools must be able to provide a Python and command-line interface (CLI) that offer significant ease of use to deep learning researchers, alongside shareable and readable configuration files that specify task definitions in a manner that can be readily ported across datasets and environments.
@@
In this work, we solve these problems with the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES offers a simple, expressive, and shareable configuration file language for task definitions and a reliable command-line- and Python-accessible tool for extracting labeled task dataframes. Task definitions in ACES are naturally separated into simple dataset-specific event predicates and dataset-agnostic inclusion/exclusion criteria, thereby permitting the same task to be used in a conceptually identical manner across diverse datasets. This approach not only enhances reproducibility but also facilitates community collaboration on task definitions, inclusion/exclusion criteria, and evaluation metrics for specific clinical use cases (Figure 1).
In contrast to prior task-definition systems, such as OHDSI’s ATLAS tool, ACES makes minimal assumptions about the input data structure or source vocabularies. In particular, ACES can be run on any dataset, provided the necessary task-specific predicates can be pre-extracted in an event-stream format111e.g., in a dataframe with columns for patient ID, timestamp, and event predicate values, and can further be run from raw data directly for any dataset in the relatively low-level and flexible Medical Event Data Standard (MEDS) [1] or EventStreamGPT (ESGPT) [10] formats. To demonstrate the utility and flexibility of ACES, we also release a collection of task definitions (Section 4) based on prior ML for healthcare works at both the dataset-agnostic criteria level and with dataset-specific predicates for the widely-used MIMIC-IV dataset [4].
@@
    ACES provides an easy-to-use utility to automatically extract these tasks from diverse sources of real-world, structured, and longitudinal electronic health record (EHR) data.
@@
Figure 1: ML task cohort extraction process with ACES and without ACES. Predicates are dataset-specific concepts that are needed to conceptually capture a machine learning task. Windows are temporal segments on a patient’s health record and are dataset-agnostic, as they are defined relative to the predicates. This distinction allows researchers to share the more complex task logic that is independent of datasets, enabling conceptual reproducibility for ML tasks in healthcare.
@@ 2 Automatic Cohort Extraction System for Event-Stream Datasets (ACES)
In this section, we introduce ACES, a novel automated task and cohort definition and extraction system that fills the key gaps in interoperability, flexibility, and accessibility left by the existing tools outlined in Section 1. To use ACES and extract a cohort for downstream ML tasks, a user only needs to do the following simple steps:
1. Install ACES: A fully functional version of ACES is pushed to PyPI, and any user can easily install it by simply running pip install es-aces. All dependencies are automatically set up with no further actions needed by the user.
2. Define Dataset: A dataset in a permitted format, such as MEDS, ESGPT, or as direct predicates, is required. More information on the data formats is available in Section 2.1.
3. Define Task: A task configuration file is required to define the task that the user wishes to extract. This configuration language is simple, clear, yet flexible, permitting users to rapidly share and iterate over task definitions for their clinical settings. Configuration specification is given in Section 2.2 and the Appendix.
4. Run the ACES CLI: ACES can be directly run from the command-line using:
@@
5. Get Outputs: The resulting output from ACES is a single unified dataframe with all valid patient instances extracted according to task specifications. Users can subsequently leverage the returned columns with original patient identifiers and health record timestamps for downstream ML tasks and benchmark creation. ACES also returns additional columns for the task label, as well as summaries of each of the predicates over windows of the patient record. Further information is available in Section 2.4 and in the example output displayed in Figure 3.
Critically, after only these five simple steps, a user can immediately, reproducibly extract a full cohort from their source dataset that matches their task definition, and begin using this task for downstream machine learning or informatics applications. Note: Due to space constraints, full technical details on all aspects of ACES and the precise details of the recursive algorithm used to extract task cohorts given a configuration file are limited to the Appendix and the online ACES documentation, available here: https://eventstreamaces.readthedocs.io/en/latest/.
@@ 2.1 Dataset Configurations
ACES is extremely flexible and can handle different data formats, including data.standard=meds [1], data.standard=esgpt [10], or data.standard=direct, where event-stream predicate views are pre-extracted by the user from any dataset schema. Note that other formats are interchangeable with these formats, such as OHDSI OMOP through the MEDS OMOP ETL [^2] https://github.com/Medical-Event-Data-Standard/meds_etl/tree/main. The use of direct predicates for extraction from datasets in a format that ACES does not natively support still significantly reduces the burden on users. Simply creating predicates is much less cumbersome than either fully converting the dataset via a CDM in order to use existing tools like ATLAS or i2b2, or performing the entire task extraction from scratch by writing in-house querying code. This demonstrates the significant improvement in utility that ACES brings across diverse data schemas compared to existing tools.
@@ 2.2 Task Configurations
In ACES, tasks are specified through configuration files that define a collection of dataset-specific event predicates, which are simple functions evaluated on individual events within a structured medical dataset. Additionally, task criteria are defined in a dataset-agnostic manner through a collection of interrelated windows, which specify regions of a patient’s record and are constrained by certain relationships. Please see Figure 2 for an example of a task configuration.
Figure 2: Example of a configuration file for the binary prediction of in-hospital mortality 48 hours after admission. References to predicates and windows are italicized and bolded, respectively. (A) Dataset-specific task predicates. These concepts are needed to conceptually capture this task and are used as constraints and boundaries for windows of the patient record. For instance, in this example, the value of “ADMISSION” denotes a hospital admission event in the external source dataset. (B) A window of the task specifying the task inputs for downstream models. Suppose we’d like to use all historic patient data up to and including 24 hours past the admission. An arbitrary criterion requiring more than 5 records can then be placed on this window to ensure that the extracted cohort contains sufficient input data (at least 5 events). (C) The trigger events for the task, which are hospital admissions as we’d like to make a mortality prediction for each admission. (D) A window of the task specifying a gap in the patient timeline. Suppose we’d like to set a minimum length of admission for our cohort (e.g., 48 hours). A temporal constraint (minimum window duration) of 48 hours could then be set to represent this requirement. (E) A window of the task specifying the task target, which is set from the end of the gap window in (D) to the immediately subsequent “discharge” or “death” predicate. This creates our binary classes for the task (i.e., $discharge=0$, $death=1$). All windows are interrelated on the patient timeline, as shown by how each window references another in the configuration file.
@@ 2.3 Command-Line Interface Arguments
The Hydra framework [21] enhances the CLI by enabling flexible run configurations and argument parsing for cohort extractions. For instance, specific arguments are required to define the external source dataset for data loading. The data.standard argument can be set to meds, esgpt, or direct, corresponding to the three supported ACES formats. Depending on the chosen format, either the path to the data file (for MEDS or direct predicates) or the path to the dataset directory (for ESGPT) must be specified to indicate the external source data from which ACES will extract the cohort. Additionally, cohort_dir and cohort_name are essential for locating and loading the task configuration file, as well as for saving results and logging operational data.
@@
For users dealing with large datasets, ACES can also be run over a collection of sharded files, extracting and storing the matching cohort for each shard individually in matching file paths. This can greatly increase computational efficiency by facilitating the processing of different shards in parallel via Hydra’s multi-run launchers333e.g., https://hydra.cc/docs/1.0/plugins/joblib_launcher/. 1\$ aces-cli \ 2 --multirun \ 3 cohort_name="<task_config_name>" \ 4 cohort_dir="/directory/to/task/config/" \ 5 data=sharded \ 6 data.standard=meds \ 7 data.root="/directory/to/dataset/shards/" \ 8 "data.shard=\\$(expand_shards <folder>/<num>)" # Sweeps over shards

### v2

@@ Abstract
Reproducibility remains a significant challenge in machine learning (ML) for healthcare. Datasets, model pipelines, and even task/cohort definitions are often private in this field, leading to a significant barrier in sharing, iterating, and understanding ML results on electronic health record (EHR) datasets. This paper addresses a significant part of this problem by introducing the Automatic Cohort Extraction System (ACES) for event-stream data. This library is designed to simultaneously simplify the development of task/cohorts for ML in healthcare and also enable the reproduction of these cohorts, both at an exact level for single datasets and at a conceptual level across datasets. To accomplish this, ACES provides (1) a highly intuitive and expressive configuration language for defining both dataset-specific concepts and dataset-agnostic inclusion/exclusion criteria, and (2) a pipeline to automatically extract patient records that meet these defined criteria from real-world data. ACES can be automatically applied to any dataset in either the Medical Event Data Standard (MEDS) or EventStreamGPT (ESGPT) formats, or to any dataset in which the necessary task-specific predicates can be extracted in an event-stream form. ACES has the potential to significantly lower the barrier to entry for defining ML tasks that learn representations, redefine the way researchers interact with EHR datasets, and significantly improve the state of reproducibility for ML studies in this modality. ACES is available at: https://github.com/justin13601/aces. A short video demonstration of ACES is available at: https://youtu.be/i_hCaHDydqA.
@@ 1 Introduction
Machine learning (ML) for healthcare suffers from a severe and systemic reproducibility crisis (McDermott et al., 2021b). This challenge is further exacerbated by the need to maintain private and secure datasets, but even with public datasets, ML pipelines are not reliably reproducible from published papers alone. For instance, in numerous attempts to reproduce ML for healthcare studies using the MIMIC-III dataset (Johnson et al., 2016), Johnson et al. found that more than half the time, the cohorts described in the studies could not be reliably reconstructed (Johnson et al., 2017).
@@
This burden in reproducing even the basic task and problem definitions in ML for healthcare studies is profoundly detrimental. Beyond the obvious concerns it raises around the robustness of reported results and their readiness for deployment, our inability to reliably define shared, canonical, reproducible task definitions limits our capacity to perform meaningful model comparisons during methodological development. This is particularly notable in settings where not all researchers have mutual access to all datasets, as is common in healthcare. Given the critical role that open benchmarks play in the advancement of ML methods (Zhang & Hardt, 2024; Salaudeen & Hardt, 2024; Shirali et al., 2023), this deficit directly translates to a significant barrier in our ability, as a research community, to effectively experiment, iterate, and develop new ML methodologies in the healthcare space.
@@
Given the clear import of this problem, the research community has naturally explored a number of prospective solutions. These can be largely categorized into two areas: (1) leveraging existing common data models (CDMs) to define reproducible task cohorts only for datasets within these schemas, and (2) defining static benchmarking tasks on individual public datasets. Both of these areas have generated numerous high-impact works. For example, in the area of CDM-driven tools, systems such as the ATLAS tool (OHDSI et al., 2024) for OHDSI’s OMOP CDM (OHDSI, 2021) and i2b2’s query tool for the i2b2 CDM (Murphy et al., 2010), as well as institution-specific tools such as the Stanford Research Repository (STARR) OMOP system and the NIHR Infections in Oxfordshire Database (IORD) cohort discovery platform have all been used to drive numerous new lines of inquiry. Unfortunately, these tools are also extremely limited in that they can only apply to the specific CDM or institutional data warehouse for which they have been defined. Further, because many of these CDMs have had limited penetrance into healthcare’s high-capacity, deep learning ecosystems, they are particularly ill-posed for task and cohort extraction within the healthcare deep learning communities. Conversely, public static benchmarks (Harutyunyan et al., 2019; McDermott et al., 2021a; van de Water et al., 2024) over datasets such as MIMIC-IV (Johnson et al., 2023) or eICU (Pollard et al., 2018) have also been extremely impactful for the ML community. However, they are all tied to only a single or small set of datasets and tasks. Given the highly dynamic nature of clinical data and healthcare requirements, this is insufficient for the benchmarking and reproducibility needs faced by the ML for healthcare community.
@@
1. The Need for Interoperability
The limited public datasets and only partially used CDMs cannot capture the diverse clinical populations, needs, and model capacities necessary for tangible ML progress in healthcare. To address this, systems for automated task extraction must be meaningfully interoperable across both public and private datasets with diverse input schemas.
2. The Need for Flexibility
A single, static benchmark cannot encompass the variety of clinical tasks relevant to clinicians and informaticians. As existing tools (with only limited interfaces to define queries using per-set vocabularies) may struggle to generalize to new clinical tasks, ideal solutions must be sufficiently flexible to accommodate a myriad of new task definitions, criteria formats, and disease or deployment areas.
3. The Need for Accessibility, Usability, & Applicability in Deep Learning Workflows
While many existing tools present with no-code interfaces (e.g., web platforms to build queries) that are essential for less technically-literate audiences, integrating such tools with deep learning workflows can prove challenging. Deep learning systems are often run in a semi-programmatic manner on siloed, private computational clusters where researchers have minimal control. Hence, existing tools can cause significant hindrance. Instead, successful software must be able to provide a Python and command-line interface (CLI) that offer significant ease of use to deep learning researchers, alongside shareable and readable configuration files that specify task definitions in a manner that can be readily ported across datasets and environments.
@@ 3. The Need for Accessibility, Usability, & Applicability in Deep Learning Workflows
In this work, we solve these problems with the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES offers a simple, expressive, and shareable configuration language for task definitions, as well as a reliable command-line interface and Python-accessible library for extracting labeled task dataframes (Figure 1).
Task definitions in ACES are naturally separated into simple dataset-specific event predicates and dataset-agnostic inclusion/exclusion criteria, thereby permitting the same task to be used in a conceptually identical manner across diverse datasets. By requiring users to specify predicates to realize their ML tasks on their specific datasets, ACES allows users to produce precise, locally specific, validatable cohorts that harmonize only the data elements needed for their task, regardless of how their input dataset is aligned or misaligned with existing ontologies or CDMs. Further, for datasets that are fully harmonized (e.g., through OHDSI vocabularies), ACES predicate definitions can be re-used across datasets without any loss of utility. In this way, ACES accommodates diverse datasets at various levels of data harmonization in a flexible, transparent manner. Overall, this approach not only enhances reproducibility but also facilitates community collaboration on task definitions, inclusion/exclusion criteria, and evaluation metrics for specific clinical use cases.
In contrast to prior task-definition systems such as ATLAS, ACES makes minimal assumptions about the input data structure or source vocabularies. In particular, ACES can be run on any dataset, provided the necessary task-specific predicates can be pre-extracted in an ”event-stream” format111e.g., in a dataframe with columns for a subject ID, timestamp, and numerical predicate values., and can further be run from raw data directly for any dataset in the relatively low-level and flexible Medical Event Data Standard (MEDS) (Arnrich et al., 2024) or EventStreamGPT (ESGPT) (McDermott et al., 2023) formats. To demonstrate the utility and flexibility of ACES, we also release a collection of task definitions (Section 4) based on prior ML for healthcare works at both the dataset-agnostic criteria level and with dataset-specific predicates for the widely-used MIMIC-IV dataset (Johnson et al., 2023).
@@
    ACES provides an easy-to-use library to automatically extract these tasks from diverse sources of real-world, structured, and longitudinal electronic health record (EHR) data.
@@
Figure 1: ML task cohort extraction process (A) with and (B) without ACES. Predicates are dataset-specific concepts that are needed to conceptually capture a machine learning task. Windows are temporal segments on a patient’s health record and are dataset-agnostic, as they are defined relative to the predicates. This distinction allows researchers to share the more complex task logic that is independent of datasets, enabling conceptual reproducibility for ML tasks in healthcare.
@@ 2 Automatic Cohort Extraction System for Event-Stream Datasets (ACES)
In this section, we introduce ACES, a novel automated task and cohort definition and extraction system that fills the key gaps in interoperability, flexibility, and accessibility left by the existing tools outlined in Section 1.
To use ACES and extract a cohort for downstream ML tasks, a user only needs to do the following simple steps:
1. Install ACES:
A fully functional version of ACES is pushed to PyPI, and any user can easily install it by simply running pip install es-aces. All dependencies are automatically set up with no further actions needed.
2. Define Dataset:
A dataset in a permitted format, such as MEDS, ESGPT, or as direct predicates, is required. More information on the data formats is available in Section 2.1.
3. Define Task:
A task configuration file is required to define the task that the user wishes to extract. This configuration language is simple, clear, yet flexible, permitting users to rapidly share and iterate over task definitions for their clinical settings. Configuration specification is given in Section 2.2 and the Appendix.
4. Run the ACES CLI:
ACES can be directly run from the command-line using:
@@ 4. Run the ACES CLI:
5. Get Outputs:
The resulting output from ACES is a single unified dataframe with all valid patient instances extracted according to task specifications. Users can subsequently leverage the returned columns with original patient identifiers and health record timestamps for downstream ML tasks and benchmark creation. ACES also returns additional columns for the task label, as well as summaries of each of the predicates over windows of the patient record. Further information is available in Section 2.4 and in the example output displayed in Figure 3.
Critically, after only these five simple steps, a user can immediately, reproducibly extract a full cohort from their source dataset that matches their task definition, and begin using this task for downstream representation learning. Due to space constraints, full technical details on all aspects of ACES and on the precise details of the recursive algorithm used to extract task cohorts given a configuration file are limited to the Appendix and the ACES public documentation, available here: https://eventstreamaces.readthedocs.io/en/latest/technical.html. Please also refer to the PDF compilation of the technical documentation in the Supplementary Material (Pages 41-51).
@@ 2.1 Dataset Configurations
ACES is extremely flexible and can handle different data formats, including data.standard=meds (Arnrich et al., 2024), data.standard=esgpt (McDermott et al., 2023), or data.standard=direct, where event-stream predicate values are pre-extracted by the user from any given dataset schema. Other CDMs are interchangeable with these formats, such as OHDSI OMOP through the MEDS OMOP ETL [^2] https://github.com/Medical-Event-Data-Standard/meds_etl/tree/main.
The use of direct predicates for extraction from datasets in a format that ACES does not natively support still significantly reduces the burden on users. Simply creating predicates is much less cumbersome than either fully converting the dataset to a CDM in order to use existing tools like ATLAS or i2b2, or performing the entire task extraction from scratch by writing in-house dataframe querying code. This demonstrates the significant improvement in utility that ACES brings across diverse data schemas compared to existing tools.
@@ 2.2 Task Configurations
In ACES, tasks are specified through configuration files that define a collection of dataset-specific event predicates, which are simple functions evaluated on individual events within a structured event-stream dataset. Additionally, task criteria are defined in a dataset-agnostic manner through a collection of interrelated windows, which specify regions of a patient’s record and are constrained by certain relationships. Please see Figure 2 for an example of a task configuration.
Figure 2: Example of a configuration file for the binary prediction of in-hospital mortality 48 hours after admission. References to predicates and windows are italicized and bolded, respectively. (A) Dataset-specific task predicates. These concepts are needed to conceptually capture this task and are used as constraints and boundaries for windows of the patient record. For instance, in this example, the value of “ADMISSION” denotes a hospital admission event in the external source dataset. (B) A window of the task specifying the task inputs for downstream models. Suppose we’d like to use all historic patient data up to and including 24 hours past the admission. An arbitrary criterion requiring more than 5 records can then be placed on this window to ensure that the extracted cohort contains sufficient input data. (C) The trigger events for the task, which are hospital admissions as we’d like to make a mortality prediction for each admission. (D) A window of the task specifying a gap in the patient timeline. Suppose we’d like to set a minimum length of admission for our cohort (e.g., 48 hours). A temporal constraint (minimum window duration) of 48 hours could then be set to represent this requirement. (E) A window of the task specifying the task target, which is set from the end of (D) to the immediately subsequent “discharge” or “death” predicate. This creates our binary classes for the task (i.e., $discharge=0$, $death=1$). All windows are interrelated on the patient timeline, as shown by how each window references another in the configuration file.
@@ 2.3 Command-Line Interface Arguments
The Hydra framework (Yadan, 2019) enhances the CLI by enabling flexible run configurations and argument parsing for cohort extractions. For instance, specific arguments are required to define the external source dataset for data loading. Depending on the chosen format (the data.standard argument), either the path to the data file (for meds or direct) or the path to the dataset directory (for esgpt) must be specified to indicate the external source data from which ACES will extract the cohort. Additionally, cohort_dir and cohort_name are essential for locating and loading the task configuration file, as well as for saving results and logging operational data.
@@
An overview of the computational profile of ACES is available in Section 4.1. Additionally, for users dealing with large datasets, ACES can also be run over a collection of sharded files, extracting and storing the matching cohort for each shard individually in matching file paths. This can greatly increase computational efficiency by facilitating the processing of different shards in parallel via Hydra’s multi-run launchers [^3] https://hydra.cc/docs/1.0/plugins/joblib_launcher/.
1\$ aces-cli \
2 --multirun \
3 cohort_name="<task_config_name>" \
4 cohort_dir="/directory/to/task/config/" \
5 data=sharded \
6 data.standard=meds \
7 data.root="/directory/to/dataset/shards/" \
8 "data.shard=\\$(expand_shards <folder>/<num>)" # Sweeps over shards